$SPAD/src/input f01qcf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)cl all
showArrayValues true
showScalarValues true
m:=5
n:=3
lda:=5
a:Matrix SF:=
 [[2.0 ,2.5 ,2.5 ],_
  [2.0 ,2.5 ,2.5 ],_
  [1.6 ,-0.4 ,2.8 ],_
  [2.0 ,-0.5 ,0.5 ],_
  [1.2 ,-0.3 ,-2.9 ]]
result:=f01qcf(m,n,lda,a,-1)

———-

References^{[1] nothing}

4